2(u-9)distributive property to romove parentheses
Found 3 solutions by peter_mccarter, miriamrios16, Alan3354:
Answer by peter_mccarter(5)

2 (u-9) distributive properties to remove parentheses
In order to remove the parentheses use the distributive property.
That is for any number a(b+c)=ab+ac
So 2(u-9)=2u-18

Answer by miriamrios16(1)

Nice picture.

Answer by Alan3354(69443)

2(u-9)
= 2*u + 2*(-9)
= 2u - 18